What should be the speed of the car at point A so that it will attain a maximum speed of 17 m/s?


1
Expert's answer
2021-04-13T06:30:49-0400

If the car moves at constant acceleration "a", and it has speed "v" at point A and speed "u=17\text{m/s}" at point B, and A and B are separated by distance "d", the speed at A can be expressed from the following equation:


"d=\frac{u^2-v^2}{2a},\\\space\\
v=\sqrt{u^2-2ad}=\sqrt{289-2ad}."